# Massless Quasiparticles in Bogoliubov-de Gennes Systems

### v1

@@ Abstract
Gapless quasiparticles can exist in the Bogoliubov-de Gennes (BdG) Hamiltonians in the mean field description of superconductors (SCs), fermionic superfluids (SFs) and quantum spin liquids (QSLs). The mechanism of gapless quasiparticles in superconductors was studied in literature based on the homotopy theory or symmetry-indicators. However, important properties of the gapless quasiparticles including the degeneracy, the energy-momentum dispersion and the responses to external probe fields need to be determined. In the present work, we investigate gapless quasiparticles in general BdG systems by using projective representation theory for the full ‘symmetry’ groups formed by combinations of lattice, spin and charge operations. We found that (I) charge conjugation (or effective charge conjugation) symmetry can yield gapless quasiparticles with linear, quadratic or higher order dispersions at high symmetry points of the Brillouin zone; (II) different quantum numbers protected level crossing can give rise to zero modes along high symmetry lines; (III) combined spatial inversion and time reversal symmetry can protect zero modes appearing at generic $k$ point. To obtain the low energy properties of gapless quasiparticles, the $k\cdot p$ theory is provided using a high efficient method–the Hamiltonian approach. Based on generalized band representation theory for BdG systems, several lattice models are constructed to illustrate the above results. Our theory provides a method to classify nodal SCs/SFs/QSLs with given symmetries, and enlightens the realization of Majorana type massless quasiparticles in condensed matter physics.
@@ I Introduction
In condensed matter physics, quasiparticle excitations emergent in the long-wave length limit behave like elementary particles in high energy physics. For instance, massless Dirac fermions or Weyl fermions can appear in semimetals at some points of the first Brillouin zone (BZ) if the crystal have certain space group symmetryWan et al. [2011], Fang et al. [2012], Wang et al. [2013], Yang and Nagaosa [2014], Weng et al. [2015], Ruan et al. [2016], Wu et al. [2020a]. These gapless quasiparticles result in observable physical consequences, such as the negative magnetic resistance effect or the existence of fermi arc spectrum in the surface excitationsNoh et al. [2017], Guo et al. [2019], Yang et al. [2019]. Furthermore, new types of quasiparticles having no counter part in high energy physics can exist in lattice systemsZyuzin and Burkov [2012], Hosur and Qi [2013], Xu et al. [2015]. The gapless quasi-particles are characterized by multiple degeneracy in the energy spectrum which carry projective representations (Reps) of their symmetry groupsYang and Liu [2017]. At a high symmetry point of the BZ, multiple degeneracy can be guaranteed by nontrivial irreducible (projective) Reps of the little co-group and around the degenerate point massless quasiparticle is formed Bradley and Davies [1968], Bradley and Cracknell [2010], Birman [2012], Hamermesh [2012]. For instance, at the K and K points of the grapheneNovoselov et al. [2005] 2-fold degeneracy is protected by a 2-dimensional (2-D) Rep of the group $\mathscr{C}_{3v}$ and Dirac-like quasiparticles are formed with linear dispersion. Gapless quasiparticles can also result from symmetry protected level crossing of two bands along certain symmetric line of the BZ, in which case the quasiparticles carry reducible (projective) Reps of the little co-groupYoung and Kane [2015], Weng et al. [2016], Bouhon and Black-Schaffer [2017], Chang et al. [2017], Yu et al. [2019], Xiao et al. [2020]. A complete description and classification of the quasiparticles needs not only the projective Reps of the symmetry groups but only the symmetry invariants which label the classes of the projective Reps Chen et al. [2013], Yang et al. [2021a], Chen et al. [2023]. With the usage of symmetry invariants, many new types of quasiparticles are found in magnetically ordered systems having weak spin-orbit coupling Wu et al. [2020b], Yang et al. [2021b], Guo et al. [2022], Tang and Wan [2022] whose symmetry groups are dubbed spin space groupsJiang et al. [2023], Xiao et al. [2023], Ren et al. [2023]. Besides the gapless quasiparticles at distinct momentum points, nodal line/surface structures can be formed if the multiple degeneracy are ensured on a line/surface of the BZ Volovik [1993], Burkov et al. [2011], Phillips and Aji [2014], Fang et al. [2015], Kim et al. [2015], Mullen et al. [2015], Bzdušek et al. [2016], Ezawa [2016], Zhao et al. [2016], Chang and Yee [2017], Yan et al. [2017]. The dispersion of the quasiparticles around the nodal points or nodal lines is determined by the effective Hamiltonian called the $k\cdot p$ model Kane [1966], Chuang and Chang [1996], Gresch [2018], Jiang et al. [2021], Yang et al. [2021c]. The $k\cdot p$ Hamiltonian, as well as the responses theory of the quasiparticle to external probe fields such as electric fields, magnetic fields and strain, can be obtained from the projective Reps associated with the quasiparticles.
@@
The quasiparticles in semimetals as discussed above have $U(1)$ charge (or particle number) conservation symmetry. Another class of quasiparticles are of Bogoliubov type when the fermions pair with each other to form Cooper pairs such that the $U(1)$ symmetry of the system breaks down to $Z_{2}$. In the BdG Hamiltonian of SCs/SFs or QSLs, the pairing between fermions yields a ‘particle-hole symmetry’ $\mathcal{P}$, which ensures that the energy spectrum is symmetric under the reflection over zero energy. Therefore, gapless Bogoliubov quasiparticles near the ‘fermi level’ (i.e. zero energy) are of special interesting science they determine the low-energy physical responses of the system. Due to the particle-hole symmetry $\mathcal{P}$, the degenerate point of the gapless quasiparticles in their energy spectrum are exactly locating at zero energy and are named zero modes (which are even-fold degenerate)Agterberg et al. [2017], Geier et al. [2020], Ono and Shiozaki [2022], Yuan et al. [2023]. For instance, $d$-wave SCs on square lattice contain zero modes at four points in the diagonal lines of the BZ, around which linearly dispersive quasiparticles arise. Especially, if charge conjugation symmetry $\mathcal{C}$ is preserved then the Bogoliubov quasiparticles are of Majorana type which are anti-particles of themselves. The ‘anti-particle’ and particle are related by charge conjugation symmetry $\mathcal{C}$ (or ‘particle-anti-particle exchange symmetry’ Bryan Gin-ge Chen [2019]). Unlike high-energy physics where Majorana fermions are massive, in condensed matter physics Majorna excitations at high symmetry points of the BZ can be massless (i.e. gapless) in $\mathcal{C}$-symmetric superconductors (SCs) or fermionic superfluids (SFs). Although Majorana zero modes (gapless chiral majorana edge states) can appear on the boundary of topological SCs/SFs, in the present work we focus on the bulk spectrum.
@@
The previous studies of bulk zero modes were mainly from the topological point of view, such as $\mathbb{Z}_{2}$ or $\mathbb{Z}$ invariants or symmetry indicators Agterberg et al. [2017], Po et al. [2017], Bradlyn et al. [2017a], Ono and Watanabe [2018], Geier et al. [2020], Ono and Shiozaki [2022]. A comprehensive theory of gapless quasiparticles in BdG systems, including their physical response to external prob fields, is still lacking. In the present work, starting from the symmetry groups and their representation theory, we systematically study the mechanisms for the appearance of the zero modes in the bulk spectrum of SCs/SFs or QSLs, and provide a method to obtain the physical properties of the gapless quasiparticles. We prove a theorem that without (effective) charge conjugation symmetry any zero modes at a given momentum point can be adiabatically removed. If (effective) charge conjugation symmetry is present, then zero modes can stably exist at high symmetry points in a band representation (with a fixed degrees of freedom in the unit cell), and gapless quasiparticles can be formed with linear, quadratic or higher order dispersions. Two types of zero modes are defined, namely irreducible zero modes and reducible zero modes, according to the projective representations of the full little co-group. From this we not only know the degeneracy of the zero modes, but can also systematically construct the $k.p$ model and the response matrix of the gapless quasiparticles to probe fields. Finally we provide a scheme to classify point nodal SCs/SFs/QSLs for a given symmetry group. We remark that our conclusions hold for all Altland–Zirnbauer symmetry classes of BdG systemsZirnbauer [1996], Altland and Zirnbauer [1997]. Furthermore, when the ‘particle-hole symmetry’ $\mathcal{P}$ is considered, the application of representations is different from that of conventional symmetry groups since $\mathcal{P}$ is anti-commuting with the Hamiltonian.
@@
The rest part of the paper is organized as follows. In section II, after a brief introduction of symmetry groups for fermionic BdG Hamiltonians, we derive the conditions for the appearance of zero modes in the bulk spectrum of SCs/SFs or QSLs. Then in section III, we provide an efficient method for obtaining the effective $k\cdot p$ model as well as the physical response to external prob fields for the gapless quasiparticles. Concrete lattice modes are provided in section IV to illustrate the results in the previous sections. Section V is devoted to a scheme for classification of point-node SCs/SFs/QSLs and a summary of the main results.
@@ II Zero modes in the Bulk spectrum of general BdG systems
In metals and insulators, the symmetry group is commuting with the Hamiltonian. For instance, at momentum point $k$ the Hamiltonian $\mathscr{H}_{k}$ is commuting with the little co-group $G_{k}$, namely $\mathscr{H}_{k}G_{k}=G_{k}\mathscr{H}_{k}$, such that the eigenspace of each energy level $\varepsilon_{k}$ generally carry an irrep of $G_{k}$. However, BdG systems like SCs/SFs are characterized by their particle-hole symmetry ${\mathcal{P}}$ which is anti-commuting with the Hamiltonian $\mathcal{P}\mathscr{H}=-\mathscr{H}\mathcal{P}$ and maps the eigenspace of $\varepsilon$ to that of $-\varepsilon$, $\mathscr{H}\mathcal{P}|\varepsilon\rangle=-\mathcal{P}\mathscr{H}|\varepsilon\rangle=-\varepsilon\mathcal{P}|\varepsilon\rangle,$ namely $\mathcal{P}|\varepsilon\rangle\propto|-\varepsilon\rangle$. Therefore, generally we need to treat the direct sum of the eigenspaces of $\{\varepsilon,-\varepsilon\}$.
@@ II.1 Symmetry operations
At mean field level, the Hamiltonian of a SC/SF/QSL reads
@@
which can be written in a matrix form in the Nambu bases $\Psi^{\dagger}=[C^{\dagger},C^{T}]$ with $C^{\dagger}=[c_{1}^{\dagger},c_{2}^{\dagger},...,c_{N}^{\dagger}]$ and $C^{T}=[c_{1},c_{2},...,c_{N}]$ ($N$ is the total number of the degrees of freedom), namely
@@
with $h_{0}={1\over 2}\sum_{i}\lambda_{i}$, $\mathscr{H}^{\dagger}=\mathscr{H}$ and ${\rm Tr}\ \mathscr{H}=0$. In the following, we will briefly introduce the symmetry operations of the Hamiltonian (1) and divide them into three different kinds.
@@
If $g$ is a spatial symmetry operation, then $gC^{\dagger}g^{-1}=C^{\dagger}u(g)$, namely, $gc_{i}^{\dagger}g^{-1}=\sum_{j=1}^{N}u_{ji}(g)c_{j}^{\dagger}$. The Hamiltonian is invariant under the action of $g$, $gHg^{-1}=H,$ or equivalently $U(g)\mathscr{H}U^{-1}(g)=\mathscr{H}$ with
@@
For a SC without magnetic order, the symmetry group is a direct product of space group and the time reversal group $Z_{2}^{\mathcal{T}}=\{E,\mathcal{T}\}$, namely the type-II magnetic space group. Here $\mathcal{T}$ acts as $\mathcal{T}C^{\dagger}\mathcal{T}^{-1}=C^{\dagger}u(\mathcal{T})K$ and $U(\mathcal{T})K\mathscr{H}KU^{-1}(\mathcal{T})=U(\mathcal{T})\mathscr{H}^{*}U^{-1}(\mathcal{T})=\mathscr{H}$ with
@@
where the three Pauli matrices $\sigma_{x,y,z}$ are the generators of the spin transformations and $\boldsymbol{n}$ is the axis of the spin rotation and $\theta\in[0,2\pi]$ is the rotation angle. If the system is free of spin-orbit coupling(SOC), then the symmetry group of the system is then a direct product group of a space group (or a magnetic space group) and $SU(2)_{\rm s}$. For systems with non-negligible SOC, the spatial point group operations also act on the spin degrees of freedom. The resulting space groups are know as double groups.
@@
Sometimes it is convenient to go back to the original Nambu bases $\Psi^{\dagger}=[C_{\uparrow}^{\dagger},C_{\downarrow}^{\dagger},C_{\uparrow}^{T},C_{\downarrow}^{T}]$, where a spin $SU(2)_{\rm s}$ rotation (2) is given by
@@
An important charge operation is the charge conjugation $\mathcal{C}$ that exchanges the creation operators with the annihilation operators, such as $\bar{\mathcal{C}}=e^{-\mathrm{i}{{\tau}_{x}\over 2}\pi}$ or $\bar{\mathcal{C}}=e^{-\mathrm{i}{{\tau}_{y}\over 2}\pi}$ in the new bases $\bar{\Psi}^{\dagger}$. Generally a charge conjugation is of order 4 with $\mathcal{C}^{2}=P_{f}$. In the original Nambu bases $\Psi^{\dagger}$ the charge conjugation takes the form of $\mathcal{C}=e^{-\mathrm{i}{\tilde{\tau}_{x}\over 2}\pi}$ or $\mathcal{C}=e^{-\mathrm{i}{\tilde{\tau}_{y}\over 2}\pi}$. A Hamiltonian is said to have charge conjugation symmetry if it is invariant under the charge conjugation transformation, namely $\hat{\mathcal{C}}H\hat{\mathcal{C}}^{-1}=\Psi^{\dagger}\mathcal{C}\mathscr{H}\mathcal{C}^{\dagger}\Psi=H,$ which requires
@@
Since the charge conjugation exchanges creation operators with annihilation operators, to preserve the $\mathcal{C}$ symmetry we must have $\lambda_{i}=0$ for all sites. Therefore, charge conjugation symmetry is a very stringent constraint in condensed matter systems since it requires the chemical potential to be zero everywhere. If a charge conjugation is combined with operations in the spin or/and lattice sectors to form a symmetry operation, then we call it an effective charge conjugation symmetry.
@@
For spin-1/2 systems with spin rotational symmetry, one can introduce the reduced Nambu bases $\psi^{\dagger}=[C_{\uparrow}^{\dagger},C_{\downarrow}^{T}]$. In the reduced Nambu bases, the charge SU(2) operations act as
$\psi^{\dagger}\to\psi^{\dagger}\exp\{-\mathrm{i}{\boldsymbol{\tau}\over 2}\cdot\boldsymbol{m}\varphi\}\otimes I_{L}$ (8)
and the particle-hole symmetry is defined as $\hat{\mathcal{P}}=\mathrm{i}\Gamma_{y}K$ with $\Gamma_{y}=\mathrm{i}\tau_{y}\otimes I_{L}$ and $\hat{\mathcal{P}}^{2}=-1$. This symmetry class is called the C-class.
@@
Since $\hat{\mathcal{P}}$ is anti-unitary, it transforms momentum $\boldsymbol{k}$ to $-\boldsymbol{k}$. Hence in momentum space, $\hat{\mathcal{P}}$ is not necessarily a symmetry operation. However, the combination of $\hat{\mathcal{P}}$ and spatial inversion (or time reversal, et al) operation form an effective particle-hole symmetry at momentum point $\boldsymbol{k}$ which anti-commute with the Hamiltonian $\mathscr{H}_{k}$. In late discussion, we denote the effective particle-hole symmetry as $\tilde{\mathcal{P}}$ with $\tilde{\mathcal{P}}=\mathcal{I}\hat{\mathcal{P}}$ (or $\tilde{\mathcal{P}}=\mathcal{T}\hat{\mathcal{P}}$, et al).
@@
A symmetry operation that commutes with the BdG Hamiltonian can be either the first kind or the second kind, or a combination of them. We write such an operation in a general form as $\alpha=\{s_{\alpha},c_{\alpha}||p_{\alpha}|t_{\alpha}\}$, where $s_{\alpha},c_{\alpha},p_{\alpha}$ respectively denote the spin, charge and lattice point group operations and $t_{\alpha}$ is a fractional translation. In later discussion, we will denote such group as $G_{f}$, the ‘fermionic’ group. When including the particle-hole ‘symmetry’, the full group will be denoted as $F_{f}=G_{f}+\mathcal{P}G_{f}$ with $\mathcal{P}^{2}=E$ or $\mathcal{P}^{2}=P_{f}$. Notice that the fermionic group $G_{f}$ is generally an extension of the space group of the lattice, and the latter is denoted as the ‘bosonic’ group $G_{b}$. For instance, in section IV.1, we will study the wallpaper group $G_{b}=p4gm\times Z_{2}^{\mathcal{T}}$ with generators $C_{4},\mathcal{T},\{\mathcal{M}_{x}|{1\over 2},{1\over 2}\}$. When ignoring the spin rotation symmetry and introducing the reduced Nambu bases, the corresponding fermionic group $G_{f}$ is generated by $C_{4}^{\prime}=(E||C_{4}),\mathcal{T}^{\prime}=(IK||\mathcal{T})$ and $\mathcal{M}_{x}^{\prime}=\{e^{-\mathrm{i}{\tau_{z}\over 2}\pi}||\mathcal{M}_{x}|{1\over 2},{1\over 2}\}$ satisfying ${\mathcal{T}^{\prime}}^{2}=1,{\mathcal{M}_{x}^{\prime}}^{2}=\{P_{f}||1,0\}$ and $G_{f}/Z_{2}^{f}=G_{b}$ (here the stands for lattice operations dressed by spin or/and charge operations) .

### v3

# Representation Theory for Massless Quasiparticles in Bogoliubov-de Gennes Systems

@@ Abstract
Linearly dispersive gapless quasiparticles can appear at general momentum points or on high symmetry momentum lines of superconductors due to topological reasons such as K theory or symmetry indicators theory. However, the zero modes associated with these quasiparticles are generally ‘accidental’ from symmetry point of view. In this work, we apply projective representation (rep) theory to analyze the bulk gapless quasiparticles in BdG systems. Different from the description of semimetals, the particle-hole ‘symmetry’ requires special treatment since it anti-commutes with the BdG Hamiltonian. Accordingly the notion of ‘simple irreducible reps (irreps)’ and ‘composite irreps’ are introduced to label the energy modes. Without charge conjugation symmetry (unitary symmetry that commutes with the Hamiltonian), no robust zero modes exist at any fixed momentum point in the bulk. However, robust zero modes at certain high-symmetry momentum points can be protected by (effective) charge conjugation symmetries, resulting in gapless quasiparticles with linear, quadratic, or higher-order dispersions. The low-energy properties of the quasiparticles, including the dispersions and responses to external probe fields, are dictated by the reps carried by these zero modes from the effective $k\cdot p$ theory. Our theory provides a framework to classify nodal superconductors/superfluids/quantum spin liquids with specified (projective) symmetry group, and sheds light on the realization of Majorana-type massless quasiparticles in condensed matter physics.
@@ I Introduction
In condensed matter physics, quasiparticle excitations emergent in the long-wavelength limit behave like elementary particles in high-energy physics. For instance, massless Dirac fermions or Weyl fermions can appear in semimetals at specific points of the first Brillouin zone (BZ) if the crystal has certain space group symmetries Wan et al. [2011], Fang et al. [2012], Wang et al. [2013], Yang and Nagaosa [2014], Weng et al. [2015], Ruan et al. [2016], Wu et al. [2020a]. These gapless quasiparticles lead to observable physical phenomena, such as negative magnetic resistance and Fermi arc spectra in surface excitations Noh et al. [2017], Guo et al. [2019], Yang et al. [2019]. Furthermore, new types of quasiparticles without counterparts in high-energy physics can exist in lattice systems Zyuzin and Burkov [2012], Hosur and Qi [2013], Xu et al. [2015]. These gapless quasiparticles are characterized by multiple degeneracies in the energy spectrum, carrying projective representations of their symmetry groups Yang and Liu [2017]. At high-symmetry points in the BZ, these degeneracies are ensured by nontrivial irreducible (projective) representations of the little co-group, forming massless quasiparticles Bradley and Davies [1968], Bradley and Cracknell [2010], Birman [2012], Hamermesh [2012]. For example, at the $K$ and $K^{\prime}$ points in graphene Novoselov et al. [2005], two-fold degeneracy is protected by a two-dimensional (2-D) representation of the group $\mathscr{C}_{3v}$, resulting in Dirac-like quasiparticles with linear dispersion. Gapless quasiparticles can also arise from symmetry-protected level crossings of two bands along certain symmetric lines in the BZ, where they carry reducible (projective) representations of the little co-group Young and Kane [2015], Weng et al. [2016], Bouhon and Black-Schaffer [2017], Chang et al. [2017], Yu et al. [2019], Xiao et al. [2020]. A comprehensive description and classification of quasiparticles require both the projective representations of symmetry groups and the symmetry invariants that label the classes of these representations Chen et al. [2013], Yang et al. [2020, 2021a], Chen et al. [2023]. With the use of symmetry invariants, many new types of quasiparticles have been discovered in magnetically ordered systems with weak spin-orbit coupling Wu et al. [2020b], Yang et al. [2021b], Guo et al. [2022], Tang and Wan [2022], whose symmetry groups are known as spin space groups Jiang et al. [2023], Xiao et al. [2023], Ren et al. [2023]. Besides isolated momentum points, nodal line and surface structures can form if multiple degeneracies are ensured along a line or surface in the BZ Volovik [1993], Burkov et al. [2011], Phillips and Aji [2014], Fang et al. [2015], Kim et al. [2015], Mullen et al. [2015], Bzdušek et al. [2016], Ezawa [2016], Zhao et al. [2016], Chang and Yee [2017], Yan et al. [2017]. The dispersion of quasiparticles around these nodal points or lines is described by the effective Hamiltonian known as the $k\cdot p$ model Kane [1966], Chuang and Chang [1996], Gresch [2018], Jiang et al. [2021], Yang et al. [2021c]. This Hamiltonian, along with the response theory of quasiparticles to external probe fields such as electric fields, magnetic fields, and strain, can be derived from the projective representations associated with the quasiparticles.
@@
The quasiparticles in semimetals discussed above have $U(1)$ charge (or particle number) conservation symmetry. Another class of quasiparticles are of Bogoliubov type, where fermions pair to form Cooper pairs, breaking the $U(1)$ symmetry down to $Z_{2}$igrist and Ueda [1991]. In the BdG Hamiltonian of SCs/SFs or QSLs, fermion pairing introduces a particle-hole symmetry $\mathcal{P}$, ensuring that the energy spectrum is symmetric around zero energy. Gapless Bogoliubov quasiparticles near the Fermi level (zero energy) are particularly interesting as they determine the system’s low-energy physical responses. Due to $\mathcal{P}$, the degenerate points of gapless quasiparticles are exactly at zero energy, termed zero modes. For instance, $d$-wave SCs on a square lattice contain zero modes at four points along the diagonal lines of the BZ, around which linearly dispersive quasiparticles arise. Notably, if charge conjugation symmetry $\mathcal{C}$ is preserved, Bogoliubov quasiparticles become Majorana types, which are their own antiparticles. The ‘antiparticle’ and particle are related by charge conjugation symmetry $\mathcal{C}$ (or ‘particle-antiparticle exchange symmetry’) Bryan Gin-ge Chen [2019]. Unlike in high-energy physics, where Majorana fermions are massive, in condensed matter physics, Majorana excitations at high-symmetry points of the BZ can be massless (gapless) in $\mathcal{C}$-symmetric SCs or SFs. Although Majorana zero modes (gapless chiral Majorana edge states) can appear on the boundary of topological SCs/SFsNyári et al. [2023], in the present work, we focus on the bulk spectrum.
@@
Previous studies of bulk zero modes have mainly focused on topological aspects, such as $\mathbb{Z}_{2}$ or $\mathbb{Z}$ invariants or symmetry indicators Slager et al. [2013], Kruthoff et al. [2017], Agterberg et al. [2017], Po et al. [2017], Bradlyn et al. [2017], Ono and Watanabe [2018], Sumita et al. [2019], Geier et al. [2020], Ono and Shiozaki [2022]. A comprehensive theory of gapless quasiparticles in BdG systems, including their physical response to external probe fields, is still lacking. In this work, starting from symmetry groups and their representation theory, we systematically study the mechanisms for the appearance of zero modes in the bulk spectrum of SCs/SFs or QSLs, and provide methods to determine the physical properties of gapless quasiparticles. We prove that without (effective) charge conjugation symmetry, zero modes at a given momentum point can be adiabatically removed. With (effective) charge conjugation symmetry, zero modes can stably exist at high-symmetry points in a band representation, forming gapless quasiparticles with linear, quadratic, or higher-order dispersions. We define two types of zero modes—irreducible and reducible—based on the projective representations of the little co-group. This allows us to determine the degeneracy of zero modes and systematically construct the $k\cdot p$ model and response matrix of gapless quasiparticles to probe fields. Finally, we propose a classification scheme for point-nodal SCs/SFs/QSLs for a given symmetry group. Our conclusions apply to all Altland–Zirnbauer symmetry classes of BdG systems.Zirnbauer [1996], Altland and Zirnbauer [1997]. Furthermore, the application of representations differs from conventional symmetry groups due to $\mathcal{P}$ anti-commuting with the Hamiltonian.
@@
The rest of the paper is organized as follows: In Section II, after a brief introduction to symmetry groups for fermionic BdG Hamiltonians, we derive the conditions for the appearance of zero modes in the bulk spectrum of SCs/SFs or QSLs. In Section III, we provide an efficient method for obtaining the effective $k\cdot p$ model and the physical response to external probe fields for gapless quasiparticles. Concrete lattice models are provided in Section IV to illustrate the results from previous sections. Section V discusses the classification scheme for point-node SCs/SFs/QSLs and summarizes the main findings.
@@ II Zero modes in the Bulk spectrum of general BdG systems
In metals and insulators, the symmetry group is commuting with the Hamiltonian. However, BdG systems like SCs/SFs are characterized by their particle-hole symmetry ${\mathcal{P}}$ which is anti-commuting with the Hamiltonian $\mathcal{P}\mathscr{H}=-\mathscr{H}\mathcal{P}$ and maps the eigenspace of $\varepsilon$ to that of $-\varepsilon$, $\mathscr{H}\mathcal{P}|\varepsilon\rangle=-\mathcal{P}\mathscr{H}|\varepsilon\rangle=-\varepsilon\mathcal{P}|\varepsilon\rangle,$ namely $\mathcal{P}|\varepsilon\rangle\propto|-\varepsilon\rangle$. Therefore, generally we need to treat the direct sum of the eigenspaces of $\{\varepsilon,-\varepsilon\}$.
@@ II.1 Symmetry operations
At the mean field level, the Hamiltonian of a SC/SF/QSL reads
@@
which can be written in matrix form in the Nambu bases $\Psi^{\dagger}=[C^{\dagger},C^{T}]$ with $C^{\dagger}=[c_{1}^{\dagger},c_{2}^{\dagger},...,c_{N}^{\dagger}]$ and $C^{T}=[c_{1},c_{2},...,c_{N}]$ ($N$ is the total number of the degrees of freedom), namely
@@
with $h_{0}={1\over 2}\sum_{i}\lambda_{i}$, $\mathscr{H}^{\dagger}=\mathscr{H}$ and ${\rm Tr}\ \mathscr{H}=0$. In the following, we will briefly introduce the symmetry operations of the Hamiltonian (1) and categorize them into three distinct types.
@@
If $g$ is a spatial symmetry operation, then
$\hat{g}C^{\dagger}\hat{g}^{-1}=C^{\dagger}u(g),$
namely, $\hat{g}c_{i}^{\dagger}\hat{g}^{-1}=\sum_{j=1}^{N}u_{ji}(g)c_{j}^{\dagger}$, where operations with $\hat{}$ acts on the Hilbert space. The Hamiltonian is invariant under the action of $g$, $\hat{g}H\hat{g}^{-1}=H,$ or equivalently $U(g)\mathscr{H}U^{-1}(g)=\mathscr{H}$ with
@@
For a SC without magnetic order, the symmetry group is a direct product of space group and the time reversal group $Z_{2}^{\mathcal{T}}=\{E,\mathcal{T}\}$, namely the type-II magnetic space group. Here $\mathcal{T}$ acts as $\hat{\mathcal{T}}C^{\dagger}\hat{\mathcal{T}}^{-1}=C^{\dagger}u(\mathcal{T})K$ and $U(\mathcal{T})K\mathscr{H}KU^{-1}(\mathcal{T})=U(\mathcal{T})\mathscr{H}^{*}U^{-1}(\mathcal{T})=\mathscr{H}$ with
@@
where the three Pauli matrices $\sigma_{x,y,z}$ are the generators of the spin transformations and $\boldsymbol{n}$ is the axis of the spin rotation and $\theta\in[0,2\pi]$ is the rotation angle. If the system is free of spin-orbit coupling(SOC), then the symmetry group of the system is then a direct product group of a space group (or a magnetic space group) and $SU(2)_{\rm s}$. For systems with non-negligible SOC, the spatial point group operations also act on the spin degrees of freedom. The resulting space groups are known as double groups.
@@
In the present work we will adopt the original Nambu bases $\Psi^{\dagger}=[C_{\uparrow}^{\dagger},C_{\downarrow}^{\dagger},C_{\uparrow}^{T},C_{\downarrow}^{T}]$, where a spin $SU(2)_{\rm s}$ rotation (2) is given by
@@
An important charge operation is the charge conjugation $\mathcal{C}$ that exchanges the creation operators with the annihilation operators, such as $\bar{\mathcal{C}}=e^{-\mathrm{i}{{\tau}_{x}\over 2}\pi}$ or $\bar{\mathcal{C}}=e^{-\mathrm{i}{{\tau}_{y}\over 2}\pi}$ in the new bases $\bar{\Psi}^{\dagger}$. Generally a charge conjugation is of order 4 with $\hat{\mathcal{C}}^{2}=\hat{P}_{f}$. In the original Nambu bases $\Psi^{\dagger}$ the charge conjugation takes the form of $\mathcal{C}=e^{-\mathrm{i}{\tilde{\tau}_{x}\over 2}\pi}$ or $\mathcal{C}=e^{-\mathrm{i}{\tilde{\tau}_{y}\over 2}\pi}$. A Hamiltonian is said to have charge conjugation symmetry if it is invariant under the charge conjugation transformation, namely $\hat{\mathcal{C}}H\hat{\mathcal{C}}^{-1}=\Psi^{\dagger}\mathcal{C}\mathscr{H}\mathcal{C}^{\dagger}\Psi=H,$ which requires
@@
Since the charge conjugation exchanges creation operators with annihilation operators, to preserve the $\mathcal{C}$ symmetry we must have $\lambda_{i}=0$ for all sites. Therefore, charge conjugation symmetry is a very stringent constraint in condensed matter systems since it requires the chemical potential to be zero everywhere. If a charge conjugation is combined with operations in the spin and/or lattice sectors to form a symmetry operation, then we call it an effective charge conjugation symmetry.
@@
For spin-1/2 systems with spin rotational symmetry, we adopt the reduced Nambu bases $\psi^{\dagger}=[C_{\uparrow}^{\dagger},C_{\downarrow}^{T}]$ in which the particle-hole symmetry is redefined as $\hat{\mathcal{P}}=\mathrm{i}\Gamma_{y}K$ with $\Gamma_{y}=\mathrm{i}\tau_{y}\otimes I_{L}$ and $\hat{\mathcal{P}}^{2}=-1$. This symmetry class is called the C-class. In the reduced Nambu bases, the $SU(2)_{s}$ group acts trivially and the $SU(2)_{c}$ operations act as
$\psi^{\dagger}\to\psi^{\dagger}\exp\{-\mathrm{i}{\boldsymbol{\tau}\over 2}\cdot\boldsymbol{m}\varphi\}\otimes I_{L}.$ (8)
@@
Since $\hat{\mathcal{P}}$ is anti-unitary, it transforms momentum $\boldsymbol{k}$ to $-\boldsymbol{k}$. Hence in momentum space, $\hat{\mathcal{P}}$ is not necessarily a symmetry operation. However, the combination of $\hat{\mathcal{P}}$ and spatial inversion (or time reversal, et al) operation form an effective particle-hole symmetry at momentum point $\boldsymbol{k}$ which anti-commute with the Hamiltonian $\mathscr{H}_{k}$. In late discussion, we denote the effective particle-hole symmetry as $\tilde{\mathcal{P}}$ with $\tilde{\mathcal{P}}=\mathcal{I}\hat{\mathcal{P}}$ (or $\tilde{\mathcal{P}}=\mathcal{T}\hat{\mathcal{P}}$, et al). Since $\{\tilde{\mathcal{P}},\mathscr{H}_{k}\}=0$, generally $\tilde{\mathcal{P}}$ and $\mathscr{H}_{k}$ cannot be diagonalized simultaneously.
A symmetry operation that commutes with the BdG Hamiltonian can be either the first kind or the second kind, or a combination of them. We write such an operation in a general form as $\alpha=\{s_{\alpha},c_{\alpha}||p_{\alpha}|t_{\alpha}\}$, where $s_{\alpha},c_{\alpha},p_{\alpha}$ respectively denote the spin, charge and lattice point group operations and $t_{\alpha}$ is a fractional translation. If the system contains spin-orbit coupling, then the spin rotations $s_{\alpha}$ are locked with the corresponding lattice point operations $p_{\alpha}$. In later discussion, we will denote the group formed by the combined operations $\alpha$ as $G_{f}=\{\alpha\}$, dubbed the ‘fermionic’ groupWang and Gu [2018], Yang et al. [2023]. Double magnetic space groups (MSGs)Bradley and Cracknell [2010] and double spin space groups (SSGs)Brinkman and Elliott [1966], Jiang et al. [2023], Xiao et al. [2023], Ren et al. [2023] are examples of $G_{f}$ which only contain one nontrivial charge operation $P_{f}$.
When including the particle-hole ‘symmetry’, the full group will be denoted as
$F_{f}=G_{f}+\mathcal{P}G_{f}$
@@
with $\mathcal{P}^{2}=E$ or $\mathcal{P}^{2}=P_{f}$. Notice that the fermionic group $G_{f}$ is generally an extension of the space group of the lattice which is denoted as the ‘bosonic’ group $G_{b}$. For instance, in section IV.1, we will study the wallpaper group $G_{b}=p4gm\times Z_{2}^{\mathcal{T}}$ with generators $C_{4},\mathcal{T},\{\mathcal{M}_{x}|{1\over 2},{1\over 2}\}$. When ignoring the spin rotation symmetry and introducing the reduced Nambu bases, the corresponding fermionic group $G_{f}$ is generated by $C_{4}^{\prime}=(E||C_{4}),\mathcal{T}^{\prime}=(IK||\mathcal{T})$ and $\mathcal{M}_{x}^{\prime}=\{e^{-\mathrm{i}{\tau_{z}\over 2}\pi}||\mathcal{M}_{x}|{1\over 2},{1\over 2}\}$ satisfying ${\mathcal{T}^{\prime}}^{2}=1,{\mathcal{M}_{x}^{\prime}}^{2}=\{P_{f}||1,0\}$ and $G_{f}/Z_{2}^{f}=G_{b}$ (here the stands for lattice operations dressed by spin or/and charge operations).